Start learning →

If P(A) = 0.35, what is the probability that A does not occur?

A0.65
B0.55
C0.35
D0.75
Answer & Solution
Correct answer: A. 0.65
Total probability is 1, so 1 - 0.35 = 0.65.
Practise this question with Quill — SSC CGL →

Keep practising Probability

Browse all Probability questions for SSC CGL, or explore more Quantitative Aptitude topics.

Start a Probability lesson →
Related questions